What is the maximum hourly output from a process with four steps having outputs of 20 units/hour, 25 units/hour, 30 units/hour and 15 units/hour? 20,25,30, or 15

Which error type indicates null hypothesis being rejected when it is true? 1,2,3,or 4

What is the performance factor of a workstation under 8-hour shift with 20 minutes cleaning and two 30-minute break and no unplanned downtime, cycle time 1 second and 18,000 items produced? .50,.75,.25,or 1.00

What is the Process Capability index (Cp) formula where USL is the Upper Specification Limit, LSL is the Lower Specification Limit and S is the standard deviation?

Question 12 options:

a)

Cp = (LSL - USL) / 3S

b)

Cp = (LSL - USL) / 6S

c)

Cp = (USL - LSL) / 3S

d)

Cp = (USL - LSL) / 6S

With 300 minutes of working time and 30 minutes of cycle time, calculate the total items which can be produced

Question 15 options:

a)

Insufficient data provided

b)

0.1

c)

9000.0

d)

10
Answer #1

1.

15

It is the lowest output of all the steps. So, it is a bottleneck and capacity of the process.

---

2.

1

It is a case of type 1 error, when a true null hypothesis is rejected.

---

4.

D

It is the formula of CP .

CP = (USL - LSL) / 6S

Here, S = Standard Deviation

---

5.

D

Total items produced = 300/30 = 10 items